Suppose the IQ of the learners in a certain University follows a normal distribution with

mean 110 and standard deviation 100?

a What proportion of the learners population have IQ's more than 90 but less than 100?

b. A random male learner is selected from this University. What is the chance that his IQ is more than 136?

c Stacy de leon, a Physics major, is prolific writer and a Math genius of this

university. If it is known that Stacy's IQ belongs to the upper 1% of the learner

population, what is her minimum IQ?


1
Expert's answer
2022-03-01T12:09:03-0500

Suppose the IQ of the learners in a certain University follows a normal distribution with mean 110 and standard deviation 10

a


"P(90<X<100)=P(Z<\dfrac{100-110}{10})"

"-P(Z\le\dfrac{90-110}{10})=P(Z<-1)-P(Z\le-2)"

"\approx0.158655-0.022750=0.1359"

b.


"P(X>136)=1-P(Z\le\dfrac{136-110}{10})"

"=1-P(Z\le\dfrac{136-110}{10})"

"=1-P(Z\le2.6)=0.004661"

c.


"P(X>x)=0.01"

"1-P(Z\le\dfrac{x-110}{10})=0.01"

"P(Z\le\dfrac{x-110}{10})=0.99"

"\dfrac{x-110}{10}=2.326348"

"x=133.26"